Show that the equation 𝑣=√t/μ



does give the



appropriate units for velocity.

1
Expert's answer
2022-02-18T15:53:27-0500

Given:


"v=\sqrt{T/\mu}"

"[v]=[N]^{1/2}[kg/m]^{-1/2}=[kg\cdot m/s^2]^{1/2}[kg/m]^{-1/2}"

"[v]=[m^2\/s^2]^{1\/2}=m\/s"